calculate the liters of a .20 M solution that can be made using 126.0 grams HNO3


1
Expert's answer
2022-03-29T13:46:04-0400

Number of moles= 126/63= 2moles

But we have 0.2 moles per litre

(2/0.2)× 1= 10L